OpenStudy (anonymous):

What do I do if there is no pi/5 on the unit circle and I converted into degrees and it gave me 36, but there are no degrees when you add or subtract that give 36

jimthompson5910 (jim_thompson5910):

What's the full problem?

jimthompson5910 (jim_thompson5910):

to evaluate cos(pi/5) ? or something?

OpenStudy (anonymous):

yea to make it a x and y @jim_thompson5910

OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by three times cosine of pi divided by five plus sine of pi divided by three times sine of pi divided by five.

jimthompson5910 (jim_thompson5910):

oh i see, so this? \[\Large \cos\left(\frac{\pi}{3}\right) * \cos\left(\frac{\pi}{5}\right) + \sin\left(\frac{\pi}{3}\right) * \sin\left(\frac{\pi}{5}\right) \]

OpenStudy (anonymous):

yes that's it but the problem is wit the pi/5

jimthompson5910 (jim_thompson5910):

you would use the identity cos(x-y) = cos(x)*cos(y) + sin(x)*sin(y)

OpenStudy (anonymous):

yea but i need to transform pi/5 into degrees and it's 36 degrees and it's not on the unit circle. I dont know how to do that

jimthompson5910 (jim_thompson5910):

no you just use cos(x-y) = cos(x)*cos(y) + sin(x)*sin(y) to convert the expression into a single trig function

jimthompson5910 (jim_thompson5910):

your answer will be of the form cos(z) where z is some number

OpenStudy (anonymous):

ohhh so it would just be cos(pi/3 - pi/5)

jimthompson5910 (jim_thompson5910):

you'd simplify that

jimthompson5910 (jim_thompson5910):

simplify pi/3 - pi/5

OpenStudy (anonymous):

okay let me try it

OpenStudy (anonymous):

Cos(2pi/15) ?

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

I was over doing it thank you :D

jimthompson5910 (jim_thompson5910):

you're welcome